2 identical satellites s1 and s2 revolve around the Earth in circular orbits at distance R and 3R respectively from the surface of the Earth,
where R is the radius of the Earth. The ratio of linear momenta of s2 to s1 is.
 
 

Grade:9

1 Answers

Adarsh
6 years ago
let the mass of the two identical satellites be m ..
on satellite s1 ,  GMm/4r2= mv2/2r   => GMm2/2r = m2v2 = p12 
on satellite s2, GMm/16r2=mv2/4r  => GMm2/4r= p22 
Now p2/p1= 1/root2 .
hope this will help you ..